OpenStudy (anonymous):

Prove the boolean identity using algebraic manipulation: WY + W'YZ' + WXZ + W'XY' = WY + W'XZ' + X'YZ' + XY'Z please helppp! :((

OpenStudy (anonymous):

i'll try. and need you join.is it ok?

OpenStudy (anonymous):

my strategy is cancel out the same term by term both sides to get the answer.

OpenStudy (anonymous):

that's ok.if you don't try,why i have to.bye

OpenStudy (anonymous):

sorry. i was away. just saw your reply

OpenStudy (anonymous):

ok. take out a paper and divide it into 2 columns, left side write out the left side of equation and right side is the right of equation, we try together. link them by = side on top

OpenStudy (anonymous):

done

OpenStudy (anonymous):

are you done? both have WY, take them out, right?

OpenStudy (anonymous):

left side W'YZ' =W'YZ'(X+X')=W'YZ'X +W'YZ'X'

OpenStudy (anonymous):

left side,too, WXZ =WXZ(Y+Y')=WXZY +WXZY' add with the first term

OpenStudy (anonymous):

do the same with all terms in 2 sides. you can see they are the same. it means the original one is equal.that's what i studied

OpenStudy (anonymous):

hope this helps

OpenStudy (anonymous):

thanks a lot...

OpenStudy (anonymous):

ok.you' re welcome